OpenStudy (anonymous):

if log_{12}27=a .calculate \[\log_{6}16 \] with using a.

OpenStudy (anonymous):

\[\log_{27}12=a \]

OpenStudy (anonymous):

can you screenshot this question, I;m not quite following it

OpenStudy (anonymous):

\[\log_{12}27=a \] \[\log_{6}16=? \]

OpenStudy (kropot72):

\[\log_{12} 27=a\] \[\log_{12} 3^{3}=a\] \[3\log_{12} 3=a\] \[\log_{12} 3=\frac{a}{3}\ ..............(1)\] \[\log_{12} 1=1\] \[\log_{12} 4=\log_{12} 12-\log_{12} 3=1-\frac{a}{3}\ .........(2)\] \[\log_{12} 16=2\log_{12} 4=2(1-\frac{a}{3})\ ..............(3)\] From equation (2): \[2\log_{12} 2=1-\frac{a}{3}\] \[\log_{12} 2=\frac{1-\frac{a}{3}}{2}\] \[\log_{12} 6=\log_{12} 3+\log_{12} 2=\frac{a}{3}+\frac{1-\frac{a}{3}}{2}\ .......(4)\] By the change of base formula: \[\log_{6} 16=\frac{\log_{12} 16}{\log_{12} 6}=\frac{2(1-\frac{a}{3})}{\frac{a}{3}+\frac{1-\frac{a}{3}}{2}}\] This simplifies to: \[\log_{6} 16=\frac{4(3-a)}{a+3}\]